OpenStudy (anonymous):

solve sqrt(20) + sqrt(45) + sqrt(12) + sqrt(27) and show all work plz

OpenStudy (anonymous):

20 = 2 * 2 * 5 = 2 sqrt(5) 45 = 3 * 3 * 5 = 3 sqrt(5) 12 = 2 * 2 * 3 = 2 sqrt(3) 27 = 3 * 3 * 3 = 3 sqrt(3) 2 sqrt(5) + 3sqrt(5) = 5 sqrt(5) 2 sqrt(3) + 3 sqrt(3) = 5 sqrt(3) 5 sqrt(5) + 5 sqrt(3)